OpenStudy (anonymous):

check my work Subtract (−4x2 + 2x) − (−5x2 + 2x3 + 3)?

OpenStudy (anonymous):

(−4x2 + 2x) − (−5x2 + 2x3 + 3) (-4x^2+2x)-(-5x^2+2x^3+3) =(-4x^2+2x)+5x^2-2x^3-3 =-4x^2+2x+5x^2-2x^3-3 =-4x^2+5x^2-2x^3+2x-3 =x^2-2x^3+2x-3 =-2x^3+x^2+2x-3 or =-(2x^3-x^2-2x+3)

OpenStudy (anonymous):

looks good to me! i wouldn't factor out a minus sign however, i would leave it at \[=-2x^3+x^2+2x-3\] but either way is correct. good work!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw you did it! i just checked.